A manufacturer is interested in the output voltage of a power supply used in a PC. Output voltage is assumed to be normally distributed, with standard deviation 0.25 Volts, and the manufacturer wishes to test using n-8 units. a) The acceptance region is 4.85 s s 5.15. Find the value of a. b) Find the power of the test for detecting a true mean output voltage of 5.1 Volts. c) Find the boundary of the critical region if the type I error probability is a-0.01 d) Find the probability of a type II error if the true mean output is 5.05 volts and a 0.01 and n-16.
Answer #1

here,

a) alpha = 1 - P(acceptance region)

= 1 - P(4.85 < X < 5.15)

we know that,

Z = (X - mean) / (standard deviation/sqrt(n)

= 1 - P((4.85-5) / (0.25/sqrt(8)) < Z < (5.15-5) / (0.25/sqrt(8)))

= 1 - P(-1.7 < Z < 1.7)

= 1 - {P(Z < 1.7) - P(Z < -1.7)}

= 1 - {0.9554-0.0446}

= 0.0892

b) calculation for the power of the test for detecting a true mean output voltage of 5.1 volts

beta = P(type II error)

= P(4.85 < X < 5.15)

= P((4.85-5.1) / (0.25/sqrt(8)) < Z < (5.15-5.1) / (0.25/sqrt(8)))

= P(-2.83 < Z < 0.57)

= 0.7157 - 0.0023

= 0.7134

so, power = 1 - beta = 1 - 0.7134 = 0.2866

c) for alpha = 0.01

Z = 2.58

critical region = mean +/- Z*standard deviation/sqrt(n)

= 5 +/- 2.58*0.25/sqrt(8)

= 5 +/- 0.228

= (4.772 , 5.228)

so the critical region would be, Xbar < 4.772 and Xbar > 5.228

d) here,

n = 16, alpha = 0.01, so , Z = 2.58

true mean output = 5.05

Xbar > 5 + Z*standard deviation/sqrt(16)

Xbar > 5 + 2.58*0.25/sqrt(16)

Xbar > 5.161

Type II error = P(X < 5.161) = P(Z < (5.161-5.05)/(0.25/sqrt(16))) = P(Z < 1.776) = 0.9621